OpenStudy (anonymous):

What is a possible first step to eliminate the variable y in the following system? 4x + 6y = 1 5x – 3y = 8

OpenStudy (anonymous):

i know i have to multiply but not sure if the first or second equation and by which number?

OpenStudy (anonymous):

my answer is multiply the second equation by 2

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

You can multiply the second one by 2, so you get a -6y term. 4x + 6y = 1 2(5x -3y) =8 ---- 4x + 6y =1 10x -6y = 8 then you can cancel the 6y's

OpenStudy (anonymous):

yep you are correct!

OpenStudy (anonymous):

thank you that is what i was thinking